# Muon tracking in a LiquidO opaque scintillator detector

### v1

@@ 2 The 64-fibre Cube detector
The skeleton (Figure 1c) is 3D printed using Elegoo ABS-Like Resin, providing a volume of 303030 mm3 to be filled with scintillator. The interior walls are made of aluminium plates which are highly reflective for the wavelengths emitted by the scintillator [30]. The top part of the Cube is enclosed using a lid (Figure 1d), which has an aluminium plate on the side facing the scintillator. Saint Gobain’s BCF-91A WLS double-clad fibres (1 mm diameter) [31] are threaded through the skeleton with a 3.2 mm spacing and sealed with a liquid insulating rubber to prevent leaks. The fibres are secured in place in the end blocks with a two-part epoxy pigmented black. The fibre ends are finished using a fly cutter, which provides a smooth end and improves photon transmission between the fibre and SiPM, which are air-coupled. The BCF-91A fibres comprise a polystyrene core surrounded by an acrylic layer and an outer fluor-acrylic layer, providing a minimum trapping efficiency of 5.6% per fibre end, as quoted in the datasheet. The WLS dye absorbs light in the 350–470 nm range, with an absorption peak at 420 nm, and emits over 460–600 nm, with an emission peak at 490 nm.
We tested both transparent and opaque scintillators in the Cube detector. The opaque scintillator used is a derivative of the wax-based opaque scintillators named “NoWaSH” [32]. The version of NoWaSH used is designed with approximately 0.5 mm scattering length [33] so that the opacity of the medium confines the majority of the light to a region on the scale of the fibre pitch. In this formulation, the NoWaSH is produced by mixing 15 wt.% wax with a base of transparent liquid scintillator, which comprises LAB, 1 wt.% PPO, and 0.03 wt.% POPOP. This LAB+PPO+POPOP transparent scintillator base is also used to acquire data, enabling a comparison between opaque and transparent scintillators. The only difference between the two scintillators is the presence of wax in the NoWaSH. The multiple fluors in the scintillator base (PPO and POPOP) allow for repeated emission/absorption to shift the scintillator photons to better match the absorption spectrum of the fibres. Figure 2 includes photos of the detector filled with transparent and opaque scintillators.
@@
The SiPM arrays used to read out the Cube and the two PixTags are Hamamatsu S13361-3050AE-08 with 64 channels in an 88 grid [34]. Each SiPM has an effective photosensitive area of 33 mm2 and a photon detection efficiency of 38% at the fibre peak emission wavelength of 490 nm. The electronics front-end modules (FEMs) plugged into each SiPM array are held in place with spring-loaded screws to the fibre end blocks to maintain consistent pressure to the face of the SiPMs (Figure 1d). The four 64-SiPM arrays used in this setup, two for the two ends of the Cube and one each for the two PixTags, are read out by the commercially available TOFPET2 ASICs developed by PETsys Electronics [35, 36]. These low-power, 64-channel chips read out and digitise signals from the SiPMs, streaming data that includes the channel number, timestamp, and charge estimate of each event. The four PETsys FEMs are connected to a PETsys FEB/D-1k module, which performs system-wide online coincidence between the FEMs and sends the data from the ASICs to the data acquisition (DAQ) computer. The FEMs are equipped with temperature sensors, allowing the regular monitoring of the temperature of the ASICs and the SiPM arrays. The temperature of the two SiPM arrays reading out the sides of the Cube were 23.5∘C and 25.0∘C, with a stability of 0.5∘C over the entire data-taking period. The entire setup was operated inside a light-tight climate chamber (see Figure 1b) with a stable air temperature of 18.00.3 ∘C.
@@ 3 Data acquisition and grouping methods
The SiPM array connected to the left (right) side of the Cube has a 51.30 V (51.90 V) breakdown voltage and is operated with 2.75 V (3.00 V) overvoltage. These voltages are configured to provide a consistent gain among the 128 SiPMs reading out the fibres, allowing the use of a single trigger threshold for all the channels. The trigger threshold is set to detect pulses generated by at least two photoelectrons (p.e.). This low threshold is selected as a trade-off between the light collection efficiency and dark noise. The SiPM arrays connected to the PixTags have a 51 V breakdown voltage and are operated with 3 V overvoltage. A trigger threshold of 5 p.e. is set in the PixTags to reduce the likelihood of noise generating a signal. By utilising the online coincidence trigger of the PETsys acquisition system, only pulses occurring in coincidence—where a signal above the threshold is detected for a PixTag and at least one other detector (either the other PixTag or either side of the Cube)—are recorded. Given that muons produce large and clean signals in the PixTags, it is likely that corresponding signals in the other detectors are caused by the same muon. Therefore, this online coincidence trigger pre-filters for cosmic muons and rejects the large majority of noise events occurring on the two sides of the Cube, where a low threshold is used. This setup gives a low overall trigger rate of 0.4 Hz, corresponding to only 4 MB of data per day.
@@ 6 Muon light cylinders
Figure 5(b) shows the distribution of the total number of p.e. collected in each muon event for the two scintillators. About 25% more light was collected when using the opaque scintillator, and furthermore, these photons were distributed across fewer SiPMs. The detected light can be estimated by considering that muons deposit approximately 2 MeV/cm and the height of the scintillator volume in the Cube is 3 cm, giving a total energy deposition of 6 MeV. In the opaque case, the average number of p.e. detected is about 170 p.e./MeV, while in the transparent is about 140 p.e./MeV. The uncertainty associated with these values of p.e./MeV is 40% and is dominated by the conversion from ToT to p.e. described in Section 5. Several factors limit the quantitative interpretation of the light levels in the two cases. For example, in the transparent setup wall reflectivity plays a more significant role in light loss than in the opaque case. Conversely, in the opaque case, the 15% of the volume that is non-scintillating wax reduces the light yield compared to transparent.
Overall, these results allow us to draw qualitative conclusions that provide insight for the optimisation of future LiquidO detector designs. Light confinement in LiquidO arises from two contributions: direct collection, in which the fibres themselves directly confine the light by stopping it from streaming away from its point of production, and stochastic confinement, given by the highly-scattering opaque medium. In the Cube detector the pitch of the fibres is 3.2 mm and their diameter is 1 mm, resulting in a relatively dense fibre lattice where the fibres occupy 7.7% of the detector volume. This configuration positions the detector within a regime where light confinement results from both direct and stochastic contributions at comparable levels. As a result, muon tracks are observed with the transparent scintillator due to the direct confinement, albeit with significantly reduced sharpness compared to the opaque case. The relative contribution of stochastic confinement can be increased by several means: using thinner fibres, increasing the fibre pitch (i.e. reducing fibre density), or employing a medium with a shorter scattering length. However, these optimisations can come with trade offs in total light collection, spatial uniformity of light collection and timing resolution of the detector response. These design choices must therefore be carefully balanced in accordance with the specific performance goals of a given LiquidO detector application.
@@ 8 Comparison with state-of-the-art muon-imaging scintillator-based detectors
LiquidO-based detectors offer clear advantages for muon imaging. The Cube prototype achieves a position resolution of 450 m at 3.2 mm pitch, 14% of the pitch, suggesting that a 15 mm-pitch LiquidO system could deliver around 2 mm resolution, compared to 2.5 mm in current triangular-bar systems. This simple scaling indicates that LiquidO has the potential to match or outperform state-of-the-art triangular-bar detectors and is expected to compete with, and potentially surpass, systems based on scintillating fibres.

### v2

@@ 2 The 64-fibre Cube detector
The skeleton (Figure 1c) is 3D printed using Elegoo ABS-Like Resin, providing a volume of 303030 mm3 to be filled with scintillator. The interior walls are made of aluminium plates which are highly reflective for the wavelengths emitted by the scintillator [30]. The top part of the Cube is enclosed using a lid (Figure 1d), which has an aluminium plate on the side facing the scintillator. Saint Gobain’s BCF-91A WLS double-clad fibres (1 mm diameter) [31] are threaded through the skeleton with a 3.2 mm spacing and sealed with a liquid insulating rubber to prevent leaks. The fibres are secured in place in the end blocks with a two-part epoxy pigmented black. The fibre ends are finished using a fly cutter, which provides a smooth end and improves photon transmission between the fibre and SiPM, which are air-coupled. The BCF-91A fibres comprise a polystyrene core surrounded by an acrylic layer and an outer fluor-acrylic layer, providing a minimum trapping efficiency of 5.6% per fibre end, as quoted in the datasheet. The WLS dye absorbs light in the 350–470 nm range, with an absorption peak at 420 nm, and emits over 460–600 nm, with an emission peak at 490 nm. The choice of 1 mm fibres was primarily guided by practical considerations: they are mechanically robust, allowing repeated filling and handling of the Cube, and they absorb scintillation photons efficiently.
We tested both transparent and opaque scintillators in the Cube detector. The opaque scintillator used is a derivative of the wax-based opaque scintillators named “NoWaSH” [32]. The version of NoWaSH used is designed with approximately 0.5 mm scattering length [33] so that the opacity of the medium confines the majority of the light to a region on the scale of the fibre pitch. In this formulation, the NoWaSH is produced by mixing 15 wt.% wax with a base of transparent liquid scintillator, which comprises LAB, 1 wt.% PPO, and 0.03 wt.% POPOP. This liquid scintillator has an absorption length on the order of meters at its peak emission wavelength and an intrinsic light yield of about 9000 photons/MeV [32]. The two fluors in the scintillator base allow for repeated emission/absorption to shift the scintillator photons to better match the absorption spectrum of the fibres. The LAB+PPO+POPOP transparent scintillator base is also used to acquire data, enabling a comparison between opaque and transparent scintillators. The only difference between the two scintillators is the presence of wax in the NoWaSH. Figure 2 includes photos of the detector filled with transparent and opaque scintillators.
@@
The SiPM arrays used to read out the Cube and the two PixTags are Hamamatsu S13361-3050AE-08 with 64 channels in an 88 grid [34]. Each SiPM has an effective photosensitive area of 33 mm2 and a photon detection efficiency of 38% at the fibre peak emission wavelength of 490 nm, as specified in the datasheet for an overvoltage of 3 V. The electronics front-end modules (FEMs) plugged into each SiPM array are held in place with spring-loaded screws to the fibre end blocks to maintain consistent pressure to the face of the SiPMs (Figure 1d). The four 64-SiPM arrays used in this setup, two for the two ends of the Cube and one each for the two PixTags, are read out by the commercially available TOFPET2 ASICs developed by PETsys Electronics [35, 36]. These low-power, 64-channel chips read out and digitise signals from the SiPMs, streaming data that includes the channel number, timestamp, and charge estimate of each event. The four PETsys FEMs are connected to a PETsys FEB/D-1k module, which performs system-wide online coincidence between the FEMs and sends the data from the ASICs to the data acquisition (DAQ) computer. The FEMs are equipped with temperature sensors, allowing the regular monitoring of the temperature of the ASICs and the SiPM arrays. The temperature of the two SiPM arrays reading out the sides of the Cube were 23.5∘C and 25.0∘C, with a stability of 0.5∘C over the entire data-taking period. The entire setup was operated inside a light-tight climate chamber (see Figure 1b) with a stable air temperature of 18.00.3 ∘C.
@@ 3 Data acquisition and grouping methods
The SiPM array connected to the left (right) side of the Cube has a breakdown voltage of 51.30 V (51.90 V) and is operated with an overvoltage of 3.50 V (3.75 V). Each array of 64 SiPMs is supplied with an independent bias voltage, allowing fine adjustment of the operating point. The small difference in overvoltage is intentionally set to equalize the gain between the two arrays, ensuring a more uniform response across all 128 SiPMs and enabling the use of a common trigger threshold for all channels. The trigger threshold is set to detect pulses generated by at least two photoelectrons (p.e.). This low threshold is selected as a trade-off between the light collection efficiency and dark noise. The SiPM arrays connected to the PixTags have a 51 V breakdown voltage and are operated with 3 V overvoltage. A trigger threshold of 5 p.e. is set in the PixTags to reduce the likelihood of noise generating a signal. By utilising the online coincidence trigger of the PETsys acquisition system, only pulses occurring in coincidence—where a signal above the threshold is detected for a PixTag and at least one other detector (either the other PixTag or either side of the Cube)—are recorded. Given that muons produce large and clean signals in the PixTags, it is likely that corresponding signals in the other detectors are caused by the same muon. Therefore, this online coincidence trigger pre-filters for cosmic muons and rejects the large majority of noise events occurring on the two sides of the Cube, where a low threshold is used. This setup gives a low overall trigger rate of 0.4 Hz, corresponding to only 4 MB of data per day.
@@ 6 Muon light cylinders
Figure 5(b) shows the distribution of the total number of p.e. collected in each muon event for the two scintillators. About 25% more light was collected when using the opaque scintillator, and furthermore, these photons were distributed across fewer SiPMs. The light level per row was reasonably uniform, with variations of approximately 10%. The detected light can be estimated by considering that muons deposit approximately 2 MeV/cm and the height of the scintillator volume in the Cube is 3 cm, giving a total energy deposition of 6 MeV. In the opaque case, the average number of p.e. detected is about 170 p.e./MeV, while in the transparent is about 140 p.e./MeV. The uncertainty associated with these values of p.e./MeV is 40% and is dominated by the conversion from ToT to p.e. described in Section 5. Several factors limit the quantitative interpretation of the light levels in the two cases. For example, in the transparent setup wall reflectivity plays a more significant role in light loss than in the opaque case. Conversely, in the opaque case, the 15% of the volume that is non-scintillating wax reduces the light yield compared to transparent.
Overall, these results allow us to draw qualitative conclusions that provide insight for the optimisation of future LiquidO detector designs. Light confinement in LiquidO arises from two contributions: direct collection, in which the fibres themselves directly confine the light by stopping it from streaming away from its point of production, and stochastic confinement, given by the highly-scattering opaque medium. In the Cube detector the pitch of the fibres is 3.2 mm and their diameter is 1 mm, resulting in a relatively dense fibre lattice where the fibres occupy 7.7% of the detector volume. This configuration positions the detector within a regime where light confinement results from both direct and stochastic contributions at comparable levels. As a result, muon tracks are observed with the transparent scintillator due to the direct confinement, albeit with significantly reduced sharpness compared to the opaque case. The relative contribution of stochastic confinement can be increased by several means: using thinner fibres, increasing the fibre pitch (i.e. reducing fibre density), or employing a medium with a shorter scattering length. However, these optimisations can come with trade offs in total light collection, spatial uniformity of light collection and timing resolution of the detector response, as well as practical considerations related to detector assembly and mechanical robustness. These design choices must therefore be carefully balanced in accordance with the specific performance goals of a given LiquidO detector application.
@@ 8 Comparison with state-of-the-art muon-imaging scintillator-based detectors
LiquidO-based detectors offer clear advantages for muon imaging. The Cube prototype achieves a position resolution of 450 m at 3.2 mm pitch, 14% of the pitch, suggesting that a 15 mm‑pitch LiquidO system could deliver around 2 mm resolution, compared to 2.5 mm in current triangular‑bar systems. This simple scaling indicates that LiquidO has the potential to match or outperform state‑of‑the‑art triangular-bar detectors and is expected to compete with, and potentially surpass, systems based on scintillating fibres.